# Propagation of Chaos for Mean-Field Langevin Dynamics and its Application to Model Ensemble

### v1

@@ Abstract
Mean-field Langevin dynamics (MFLD) is an optimization method derived by taking the mean-field limit of noisy gradient descent for two-layer neural networks in the mean-field regime. Recently, the propagation of chaos (PoC) for MFLD has gained attention as it provides a quantitative characterization of the optimization complexity in terms of the number of particles and iterations. A remarkable progress by Chen et al. (2022) showed that the approximation error due to finite particles remains uniform in time and diminishes as the number of particles increases. In this paper, by refining the defective log-Sobolev inequality—a key result from that earlier work—under the neural network training setting, we establish an improved PoC result for MFLD, which removes the exponential dependence on the regularization coefficient from the particle approximation term of the optimization complexity. As an application, we propose a PoC-based model ensemble strategy with theoretical guarantees.
@@ Introduction
A two layer mean-field neural network (MFNN) with $N$ neurons is defined as an empirical average of $N$ functions: $\mathbb{E}_{X\sim\rho_{\mathbf{x}}}\left[h\left(X,\cdot\right)\right]=\frac{1}{N}\sum^{N}_{i=1}h(x^{i},\cdot)$, where each $h(x^{i},\cdot)$ represents a single neuron with parameter $x^{i}$ and $\rho_{\mathbf{x}}=\frac{1}{N}\sum^{N}_{i=1}\delta_{x_{i}}$ is an empirical distribution. As the number of neurons get infinitely large $(N\rightarrow\infty)$, the mean-field limit is attained: $\rho_{\mathbf{x}}\rightarrow\mu$, leading to MFNN having an infinite number of particles: $\mathbb{E}_{X\sim\mu}\left[h(X,\cdot)\right]$. Since a distribution $\mu$ parameterizes the model in this mean-field limit, training can now be formulated as the optimization over the space of probability distributions (Nitanda & Suzuki, 2017). Gradient descent for MFNNs exhibits global convergence (Chizat & Bach, 2018; Mei et al., 2018) and adaptivity (Yang & Hu, 2020; Ba et al., 2022). To improve stability during training, one may consider noisy gradient training by adding Gaussian noise, giving rise to mean-field Langevin dynamics (MFLD) (Mei et al., 2018; Hu et al., 2019). MFLD, with $N=\infty$, also achieves global convergence to the optimal solution (Hu et al., 2019; Jabir et al., 2019), with an exponential convergence rate under the uniform log-Sobolev inequality (LSI) Nitanda et al. (2022); Chizat (2022) in the continuous-time setting.
However, the mean-field limit attained at $N=\infty$ cannot be accurately replicated in real-life scenarios. When employing a finite-particle system $\rho_{\mathbf{x}}$, the approximation error that arises has been studied in the literature on propagation of chaos (PoC) Sznitman (1991). In the context of MFLD, Chen et al. (2022); Suzuki et al. (2023a) proved the uniform-in-time PoC for the trajectory of MFLD. In particular, in the long-time limit, they established the bounds $\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})=O\left(\frac{\lambda}{\alpha N}\right)$, where $\alpha\gtrsim\exp\left(-\Theta\left(\frac{1}{\lambda}\right)\right)$ is the LSI constant on proximal Gibbs distributions, $\lambda$ is the regularization coefficient, and $\mathcal{L}^{(N)}(\mu^{(N)}_{*})$ and $\mathcal{L}(\mu_{*})$ are the optimal values in finite- and infinite-particle systems. Subsequently, Nitanda (2024) improved upon this result by removing $\alpha$ from the above bound, resulting in $O\left(\frac{1}{N}\right)$. This refinement of the bound is significant as previously, the LSI constant could become exponentially small as $\lambda\rightarrow 0$. While Nitanda (2024) also established PoC for the MFLD trajectory by incorporating the uniform-in-N LSI Chewi et al. (2024): $\mathcal{L}^{(N)}(\mu^{(N)}_{t})\to\mathcal{L}^{(N)}(\mu^{(N)}_{*})$, this approach is indirect for showing convergence to the mean-field limit $\mathcal{L}(\mu_{*})$ and results in a slower convergence rate over time.
@@ 1.1 Contributions
The PoC for MFLD Chen et al. (2022); Suzuki et al. (2023a) consists of particle approximation error $O\left(\frac{\lambda}{\alpha N}\right)$ due to finite-$N$-particles and optimization error $\exp(-\Theta(\lambda\alpha t))$. This result basically builds upon the defective LSI: $\exists\delta>0$,
@@
implicitly established by Chen et al. (2022) under the uniform LSI condition Nitanda et al. (2022); Chizat (2022), where $\mathrm{FI}$ is Fisher information. The dependence on LSI-constant $\alpha$ in $O\left(\frac{\lambda}{\alpha N}\right)$ of PoC is basically inherited from $\delta$. In our work, we first remove the dependence on $\alpha$ from $\delta$ by introducing uniform directional LSI (Assumption 2) in training MFNNs setting. Based on this defective LSI, we then derive an improved PoC for MFLD where the particle approximation error is $O\left(\frac{1}{N}\right)$. Similar to Nitanda (2024), this improvement exponentially reduces the required number of particles since the constant $\alpha\gtrsim\exp\left(-\Theta(\frac{1}{\lambda})\right)$ can exponentially decrease as $\lambda\to\infty$. Moreover, our result demonstrates a faster optimization speed compared to Nitanda (2024); Chewi et al. (2024) due to a different exponent $\alpha$ in the optimization error terms: $\exp(-\Theta(\lambda\alpha t))$. In our analysis, $\alpha$ is a constant of the uniform directional LSI, which is larger than the LSI constant on $\mu^{(N)}_{*}$ appearing in the optimization error in Nitanda (2024); Chewi et al. (2024) (see the discussion following Theorem 1).
@@
Next, we translate the PoC result regarding objective gap into the point-wise and uniform model approximation errors: $|\mathbb{E}_{X\sim\rho_{\mathbf{x}}}[h(X,z)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]|$ and $\|\mathbb{E}_{X\sim\rho_{\mathbf{x}}}[h(X,\cdot)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,\cdot)]\|_{\infty}$ useful for obtaining generalization error bound on classification task Suzuki et al. (2023b); Nitanda et al. (2024). Again, the bound consists of the sum of particle approximation and optimization error terms. Compared to the previous results Suzuki et al. (2023a, b), our bound is tighter since the particle approximation term is independent of the LSI-constant. This improvement directly eliminates the requirement for an exponential number of neurons with respect to dimension $d$ in their learning setup (e.g., $k$-parity problems Suzuki et al. (2023b)). We also propose a PoC-based model ensemble method to further reduce the model approximation error and empirically verify its performance on synthetic datasets. To our knowledge, our study is the first to provide a theoretical guarantee for model ensembling of MFNNs using PoC results. Going beyond the scope of the theory, we examine the applicability of our method to merging LoRA parameters for language models.
@@
  • We demonstrate an improved PoC for MFLD (Theorem 1) under uniform directional LSI condition (Assumption 2). This improvement removes the dependence on LSI constant $\alpha\gtrsim\exp\left(-\Theta\left(\frac{1}{\lambda}\right)\right)$ from the particle approximation error in Chen et al. (2022); Suzuki et al. (2023a) and accelerates the optimization speed in Nitanda (2024); Chewi et al. (2024).
@@
  • We translate the PoC result regarding objective gap into point-wise and uniform model approximation errors (Theorems 2, 3, and 4). These results also remove the dependence on the LSI constant from the particle approximation terms in the previous model approximation errors Suzuki et al. (2023a, b).
@@ 2.1 Problem setting
where $r:\mathbb{R}^{d}\to\mathbb{R}$ is a $\lambda^{\prime}$-strongly convex function (e.g., $r(x)=\lambda^{\prime}\|x\|_{2}^{2}$ $(\lambda^{\prime}>0)$). We set $F(\mu)=F_{0}(\mu)+\mathbb{E}_{\mu}[r(X)]$. A typical example of $F_{0}$ is an empirical risk of the two-layer mean-field neural network (see Example 1). Throughout the paper, we assume that the solution $\mu_{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ of the problem (2) exists and make the following regularity assumption Chizat (2022); Nitanda et al. (2022); Chen et al. (2023) under which $\mu_{*}$ is unique and satisfies the optimality condition: $\mu_{*}\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F(\mu_{*})}{\delta\mu}\right)$ (see Hu et al. (2019); Chizat (2022) for the details).
@@ 2.2 Mean-field Langevin dynamics and uniform-in-time propagation of chaos
where $\{W_{t}^{i}\}_{t\geq 0},~{}(i\in\{1,\ldots,N\})$ are independent standard Brownian motions. Note that Eq. (4) is equivalent to the Langevin dynamics $\mathrm{d}\mathbf{X}_{t}=-N\nabla_{\mathbf{X}}F(\rho_{\mathbf{X}_{t}})\mathrm{d}t+\sqrt{2\lambda}\mathrm{d}\mathbf{W}_{t}$ on $\mathbb{R}^{dN}$, where $\{\mathbf{W}_{t}\}_{t\geq 0}$ is the standard Brownian motion on $\mathbb{R}^{dN}$ since $N\nabla_{x^{i}}F(\rho_{\mathbf{x}})=\nabla\frac{\delta F(\mu_{\mathbf{x}})}{\delta\mu}(x^{i})$ (Chizat, 2022). Therefore, $\mu^{(N)}_{t}=\mathrm{Law}(\mathbf{X}_{t})$ converges to the Gibbs distribution
@@
where $\{W_{t}\}_{t\geq 0}$ is the $d$-dimensional standard Brownian motion with $W_{0}=0$. Under the log-Sobolev inequality on the proximal Gibbs distribution $\hat{\mu}\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F}{\delta\mu}\right)$, Nitanda et al. (2022); Chizat (2022) showed the exponential convergence of the objective gap $\mathcal{L}(\mu_{t})-\mathcal{L}(\mu_{*})$, where $\mu_{*}=\mathop{\mathrm{argmin}}_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\mathcal{L}(\mu)$.
@@
This approximation error has been studied in the literature of PoC. Recently, Suzuki et al. (2023a) proved the following uniform-in-time PoC for Eq. (3) by using the techniques in Chen et al. (2022):
@@
where $\Delta^{(N)}_{0}=\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\mathcal{L}(\mu_{*})$ is the initial gap and $\delta_{\eta,N}=\frac{(\lambda\eta+\eta^{2})D_{1}}{\lambda\alpha}+\frac{\lambda D_{2}}{\alpha N}$ $(\exists D_{1},D_{2}>0)$ is the discretization error in time and space. The continuous-time counterpart ($\eta\to 0$) was proved by Chen et al. (2022). The typical estimation of LSI-constant $\alpha\gtrsim\exp(-\Theta(1/\lambda))$ (e.g., Theorem 1 in Suzuki et al. (2023a)) using Holley and Stroock argument (Holley & Stroock, 1987) or Miclo’s trick (Bardet et al., 2018)) suggests the exponential blow-up of the particle approximation error $\frac{\lambda D_{2}}{\alpha N}$ in Eq. (7) as $\lambda\to 0$.
Afterward, this exponential dependence was removed by Nitanda (2024); Chewi et al. (2024) that evaluate the particle approximation error at the solution: $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})$ and optimization error: $\frac{1}{N}\left(\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right)$, respectively. In the risk minimization problem setting, Nitanda (2024) proved $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{C}{N}$ $(\exists C>0)$ and Chewi et al. (2024) proved uniform-in-$N$ LSI on $\mu^{(N)}_{*}\in\mathbb{R}^{dN}$ with the constant estimation $\bar{\alpha}\gtrsim\frac{\lambda^{\prime}}{\lambda}\exp\left(-O\left(\frac{1}{\lambda^{\prime}}+\frac{1}{\lambda\lambda^{\prime}}+\frac{1}{\lambda^{2}\lambda^{\prime 3}}\right)\right)$, leading to the $N$-independent convergence rate of $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}(\mu_{*})$ up to the particle approximation error $C/N$ plus time-discretization error.
@@
The nonlinearity of $F_{0}$ is the key to the PoC analysis for mean-field models, thereby motivating the use of the Bregman divergence associated with $F_{0}$ (Nitanda, 2024); for distributions $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
We give an example of training MFNNs, which satisfies Assumptions 2, 3, and 4.

### v2

@@ Abstract
Mean-field Langevin dynamics (MFLD) is an optimization method derived by taking the mean-field limit of noisy gradient descent for two-layer neural networks in the mean-field regime. Recently, the propagation of chaos (PoC) for MFLD has gained attention as it provides a quantitative characterization of the optimization complexity in terms of the number of particles and iterations. A remarkable progress by Chen et al., (2022) showed that the approximation error due to finite particles remains uniform in time and diminishes as the number of particles increases. In this paper, by refining the defective log-Sobolev inequality—a key result from that earlier work—under the neural network training setting, we establish an improved PoC result for MFLD, which removes the exponential dependence on the regularization coefficient from the particle approximation term of the optimization complexity. As an application, we propose a PoC-based model ensemble strategy with theoretical guarantees.
@@ Introduction
A two layer mean-field neural network (MFNN) with $N$ neurons is defined as an empirical average of $N$ functions: $\mathbb{E}_{X\sim\rho_{\mathbf{x}}}\left[h\left(X,\cdot\right)\right]=\frac{1}{N}\sum^{N}_{i=1}h(x^{i},\cdot)$, where each $h(x^{i},\cdot)$ represents a single neuron with parameter $x^{i}$ and $\rho_{\mathbf{x}}=\frac{1}{N}\sum^{N}_{i=1}\delta_{x_{i}}$ is an empirical distribution. As the number of neurons get infinitely large $(N\rightarrow\infty)$, the mean-field limit is attained: $\rho_{\mathbf{x}}\rightarrow\mu$, leading to MFNN having an infinite number of particles: $\mathbb{E}_{X\sim\mu}\left[h(X,\cdot)\right]$. Since a distribution $\mu$ parameterizes the model in this mean-field limit, training can now be formulated as the optimization over the space of probability distributions (Nitanda and Suzuki,, 2017). Gradient descent for MFNNs exhibits global convergence (Chizat and Bach,, 2018; Mei et al.,, 2018) and adaptivity (Yang and Hu,, 2020; Ba et al.,, 2022). To improve stability during training, one may consider noisy gradient training by adding Gaussian noise, giving rise to mean-field Langevin dynamics (MFLD) (Mei et al.,, 2018; Hu et al.,, 2019). MFLD, with $N=\infty$, also achieves global convergence to the optimal solution (Hu et al.,, 2019; Jabir et al.,, 2019), with an exponential convergence rate under the uniform log-Sobolev inequality (LSI) Nitanda et al., (2022); Chizat, (2022) in the continuous-time setting.
However, the mean-field limit attained at $N=\infty$ cannot be accurately replicated in real-life scenarios. When employing a finite-particle system $\rho_{\mathbf{x}}$, the approximation error that arises has been studied in the literature on propagation of chaos (PoC) Sznitman, (1991). In the context of MFLD, Chen et al., (2022); Suzuki et al., 2023a proved the uniform-in-time PoC for the trajectory of MFLD. In particular, in the long-time limit, they established the bounds $\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})=O\left(\frac{\lambda}{\alpha N}\right)$, where $\alpha\gtrsim\exp\left(-\Theta\left(\frac{1}{\lambda}\right)\right)$ is the LSI constant on proximal Gibbs distributions, $\lambda$ is the regularization coefficient, and $\mathcal{L}^{(N)}(\mu^{(N)}_{*})$ and $\mathcal{L}(\mu_{*})$ are the optimal values in finite- and infinite-particle systems. Subsequently, Nitanda, (2024) improved upon this result by removing $\alpha$ from the above bound, resulting in $O\left(\frac{1}{N}\right)$. This refinement of the bound is significant as previously, the LSI constant could become exponentially small as $\lambda\rightarrow 0$. While Nitanda, (2024) also established PoC for the MFLD trajectory by incorporating the uniform-in-N LSI Chewi et al., (2024): $\mathcal{L}^{(N)}(\mu^{(N)}_{t})\to\mathcal{L}^{(N)}(\mu^{(N)}_{*})$, this approach is indirect for showing convergence to the mean-field limit $\mathcal{L}(\mu_{*})$ and results in a slower convergence rate over time.
@@ 1.1 Contributions
The PoC for MFLD Chen et al., (2022); Suzuki et al., 2023a consists of particle approximation error $O\left(\frac{\lambda}{\alpha N}\right)$ due to finite-$N$-particles and optimization error $\exp(-\Theta(\lambda\alpha t))$. This result basically builds upon the defective LSI: $\exists\delta>0$,
@@
implicitly established by Chen et al., (2022) under the uniform LSI condition Nitanda et al., (2022); Chizat, (2022), where $\mathrm{FI}$ is Fisher information. The dependence on LSI-constant $\alpha$ in $O\left(\frac{\lambda}{\alpha N}\right)$ of PoC is basically inherited from $\delta$. In our work, we first remove the dependence on $\alpha$ from $\delta$ by introducing uniform directional LSI (Assumption 2) in training MFNNs setting. Based on this defective LSI, we then derive an improved PoC for MFLD where the particle approximation error is $O\left(\frac{1}{N}\right)$. Similar to Nitanda, (2024), this improvement exponentially reduces the required number of particles since the constant $\alpha\gtrsim\exp\left(-\Theta(\frac{1}{\lambda})\right)$ can exponentially decrease as $\lambda\to\infty$. Moreover, our result demonstrates a faster optimization speed compared to Nitanda, (2024); Chewi et al., (2024) due to a different exponent $\alpha$ in the optimization error terms: $\exp(-\Theta(\lambda\alpha t))$. In our analysis, $\alpha$ is a constant of the uniform directional LSI, which is larger than the LSI constant on $\mu^{(N)}_{*}$ appearing in the optimization error in Nitanda, (2024); Chewi et al., (2024) (see the discussion following Theorem 1).
@@
Next, we translate the PoC result regarding objective gap into the point-wise and uniform model approximation errors: $|\mathbb{E}_{X\sim\rho_{\mathbf{x}}}[h(X,z)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,z)]|$ and $\|\mathbb{E}_{X\sim\rho_{\mathbf{x}}}[h(X,\cdot)]-\mathbb{E}_{X\sim\mu_{*}}[h(X,\cdot)]\|_{\infty}$ useful for obtaining generalization error bound on classification task Suzuki et al., 2023b ; Nitanda et al., (2024). Again, the bound consists of the sum of particle approximation and optimization error terms. Compared to the previous results Suzuki et al., 2023a ; Suzuki et al., 2023b , our bound is tighter since the particle approximation term is independent of the LSI-constant. This improvement directly eliminates the requirement for an exponential number of neurons with respect to dimension $d$ in their learning setup (e.g., $k$-parity problems Suzuki et al., 2023b ). We also propose a PoC-based model ensemble method to further reduce the model approximation error and empirically verify its performance on synthetic datasets. To our knowledge, our study is the first to provide a theoretical guarantee for model ensembling of MFNNs using PoC results. Going beyond the scope of the theory, we examine the applicability of our method to merging LoRA parameters for language models.
@@
  • We demonstrate an improved PoC for MFLD (Theorem 1) under uniform directional LSI condition (Assumption 2). This improvement removes the dependence on LSI constant $\alpha\gtrsim\exp\left(-\Theta\left(\frac{1}{\lambda}\right)\right)$ from the particle approximation error in Chen et al., (2022); Suzuki et al., 2023a and accelerates the optimization speed in Nitanda, (2024); Chewi et al., (2024).
@@
  • We translate the PoC result regarding objective gap into point-wise and uniform model approximation errors (Theorems 2, 3, and 4). These results also remove the dependence on the LSI constant from the particle approximation terms in the previous model approximation errors Suzuki et al., 2023a ; Suzuki et al., 2023b .
@@ 2.1 Problem setting
where $r:\mathbb{R}^{d}\to\mathbb{R}$ is a $\lambda^{\prime}$-strongly convex function (e.g., $r(x)=\lambda^{\prime}\|x\|_{2}^{2}$ $(\lambda^{\prime}>0)$). We set $F(\mu)=F_{0}(\mu)+\mathbb{E}_{\mu}[r(X)]$. A typical example of $F_{0}$ is an empirical risk of the two-layer mean-field neural network (see Example 1). Throughout the paper, we assume that the solution $\mu_{*}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ of the problem (2) exists and make the following regularity assumption Chizat, (2022); Nitanda et al., (2022); Chen et al., (2023) under which $\mu_{*}$ is unique and satisfies the optimality condition: $\mu_{*}\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F(\mu_{*})}{\delta\mu}\right)$ (see Hu et al., (2019); Chizat, (2022) for the details).
@@ 2.2 Mean-field Langevin dynamics and uniform-in-time propagation of chaos
where $\{W_{t}^{i}\}_{t\geq 0},~(i\in\{1,\ldots,N\})$ are independent standard Brownian motions. Note that Eq. (4) is equivalent to the Langevin dynamics $\mathrm{d}\mathbf{X}_{t}=-N\nabla_{\mathbf{X}}F(\rho_{\mathbf{X}_{t}})\mathrm{d}t+\sqrt{2\lambda}\mathrm{d}\mathbf{W}_{t}$ on $\mathbb{R}^{dN}$, where $\{\mathbf{W}_{t}\}_{t\geq 0}$ is the standard Brownian motion on $\mathbb{R}^{dN}$ since $N\nabla_{x^{i}}F(\rho_{\mathbf{x}})=\nabla\frac{\delta F(\mu_{\mathbf{x}})}{\delta\mu}(x^{i})$ (Chizat,, 2022). Therefore, $\mu^{(N)}_{t}=\mathrm{Law}(\mathbf{X}_{t})$ converges to the Gibbs distribution
@@
where $\{W_{t}\}_{t\geq 0}$ is the $d$-dimensional standard Brownian motion with $W_{0}=0$. Under the log-Sobolev inequality on the proximal Gibbs distribution $\hat{\mu}\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F(\mu)}{\delta\mu}\right)$, Nitanda et al., (2022); Chizat, (2022) showed the exponential convergence of the objective gap $\mathcal{L}(\mu_{t})-\mathcal{L}(\mu_{*})$, where $\mu_{*}=\mathop{\mathrm{argmin}}_{\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})}\mathcal{L}(\mu)$.
@@
This approximation error has been studied in the literature of PoC. Recently, Suzuki et al., 2023a proved the following uniform-in-time PoC for Eq. (3) by using the techniques in Chen et al., (2022):
@@
where $\Delta^{(N)}_{0}=\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{0})-\mathcal{L}(\mu_{*})$ is the initial gap and $\delta_{\eta,N}=\frac{(\lambda\eta+\eta^{2})D_{1}}{\lambda\alpha}+\frac{\lambda D_{2}}{\alpha N}$ $(\exists D_{1},D_{2}>0)$ is the discretization error in time and space. The continuous-time counterpart ($\eta\to 0$) was proved by Chen et al., (2022). The typical estimation of LSI-constant $\alpha\gtrsim\exp(-\Theta(1/\lambda))$ (e.g., Theorem 1 in Suzuki et al., 2023a ) using Holley and Stroock argument (Holley and Stroock,, 1987) or Miclo’s trick (Bardet et al.,, 2018)) suggests the exponential blow-up of the particle approximation error $\frac{\lambda D_{2}}{\alpha N}$ in Eq. (7) as $\lambda\to 0$.
Afterward, this exponential dependence was removed by Nitanda, (2024); Chewi et al., (2024) that evaluate the particle approximation error at the solution: $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})$ and optimization error: $\frac{1}{N}\left(\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}^{(N)}(\mu^{(N)}_{*})\right)$, respectively. In the risk minimization problem setting, Nitanda, (2024) proved $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{*})-\mathcal{L}(\mu_{*})\leq\frac{C}{N}$ $(\exists C>0)$ and Chewi et al., (2024) proved uniform-in-$N$ LSI on $\mu^{(N)}_{*}\in\mathbb{R}^{dN}$ with the constant estimation $\bar{\alpha}\gtrsim\frac{\lambda^{\prime}}{\lambda}\exp\left(-O\left(\frac{1}{\lambda^{\prime}}+\frac{1}{\lambda\lambda^{\prime}}+\frac{1}{\lambda^{2}\lambda^{\prime 3}}\right)\right)$, leading to the $N$-independent convergence rate of $\frac{1}{N}\mathcal{L}^{(N)}(\mu^{(N)}_{k})-\mathcal{L}(\mu_{*})$ up to the particle approximation error $C/N$ plus time-discretization error.
@@
Remark.
Wang, (2024) also introduced the conditional Gibbs distribution and imposed a Poincaré inequality on it.
@@
The nonlinearity of $F_{0}$ is the key to the PoC analysis for mean-field models, thereby motivating the use of the Bregman divergence associated with $F_{0}$ (Nitanda,, 2024); for distributions $\mu,\mu^{\prime}\in\mathcal{P}_{2}(\mathbb{R}^{d})$,
@@
Here, we give a connection between a conditional Gibbs distribution $\nu_{i|-i}(\cdot|\mathbf{x}^{-i})$ and proximal Gibbs distribution $\hat{\rho}_{\mathbf{x}}$ using the Bregman divergence. Given $\mathbf{x}=(x^{1},\ldots,x^{N})$, the following relationship holds as a probability distribution over $x\in\mathbb{R}$:
$\frac{\mathrm{d}\nu_{i|-i}}{\mathrm{d}x}$ $(x|\mathbf{x}^{-i})\propto\exp\left(-\frac{N}{\lambda}F(\rho_{x\cup\mathbf{x}^{-i}})\right)$
$=\exp\biggl{(}-\frac{N}{\lambda}\biggl{(}F(\rho_{\mathbf{x}})+\left\langle\frac{\delta F(\rho_{\mathbf{x}})}{\delta\mu},\rho_{x\cup\mathbf{x}^{-i}}-\rho_{\mathbf{x}}\right\rangle$
$+B_{F}(\rho_{x\cup\mathbf{x}^{-i}},\rho_{\mathbf{x}})\biggr{)}\biggr{)}$
$\propto\exp\left(-\frac{1}{\lambda}\frac{\delta F(\rho_{\mathbf{x}})}{\delta\mu}(x)-\frac{N}{\lambda}B_{F}(\rho_{x\cup\mathbf{x}^{-i}},\rho_{\mathbf{x}})\right)$
$\propto\frac{\mathrm{d}\hat{\rho}_{\mathbf{x}}}{\mathrm{d}x}(x)\exp\left(-\frac{N}{\lambda}B_{F}(\rho_{x\cup\mathbf{x}^{-i}},\rho_{\mathbf{x}})\right).$ (8)
This connection is useful for deriving the LSI on $\nu_{i|-i}$ from the uniform LSI on the proximal Gibbs distributions $\hat{\rho}_{\mathbf{x}}$ (see Nitanda et al., (2022)), in combination with Assumption 4 and Holley-Strook argument.
We give an example of training MFNNs that satisfies Assumptions 2, 3, and 4.